Night Report 2016-02-17

Observing: Stefan Taubenberger, Tamar Faran, Markus Kromer

Support Team: Nancy Elias-Rosa

TAT: Morgan Fraser, Lluís Galbany

Weather conditions

Start of night: clear, humidity ~35%, wind ~1.5 m/s

Middle of night: clear, humidity ~35%, wind ~ 3m/s

End of night: clear, humidity ~40%, wind ~ 5m/s




Morning: -


00:30 UT: data analysis

00:50 UT: Classification: Gaia16ady, gr13, 480s, seeing ~1.3"

01:05 UT: Follow-up: ASASSN-15pz, gr13, 2x2700s, seeing ~1.2"

02:50 UT: Standard: L745a, gr13, 100s, seeing ~1.1"

03:00 UT: Classification: LSQ16pz, gr13, 2700s, seeing ~1.1"

03:50 UT: Classification: PS16aia, source faded

04:00 UT: Follow-up: SN2016B, gr13, 900s, seeing ~1"

04:20 UT: Classification: ATLAS16aas, gr13, 900s, seeing ~0.9"


05:00 UT: Follow-up: SN2015F, K, 1500s, seeing ~0.7"

06:10 UT: Follow-up: SN2016B, BG + RG, 960s + 1500s, seeing ~0.9"

07:40 UT: Telluric standard: HIP60074, BG + RG, 10s, seeing ~1.2"

07:45 UT: Follow-up: SN2016adj, JHK + BG + RG, 40s + 60s + 120s, seeing ~0.7"

09:00 UT: Telluric standard: HIP64550, BG + RG, 4.5s + 6s, seeing ~1.2"

09:05 UT: Standard: EG274, BG + RG, 60s + 120s, seeing ~0.9"

09:20 UT: End of the night

* Towards the end of the night there were some instrumental problems with SOFI - the software did not read the right exposure time from the OB's, and it was increased by around a factor of 2. We also lost some observation time due to a problem with the count down, which for some reason did not stop at 0:00 and went into negative values of almost 3 minutes. We hope that these problems did not affect the images.